OpenStudy (anonymous):

how do we solve the initial-value problem? −7y''−7y=sin(x); y(0)=2 y'(0)=−4

OpenStudy (anonymous):

are you all in systems of first order eq yet?

OpenStudy (anonymous):

yea we passed that section@

OpenStudy (anonymous):

do you remember how to set up the matrix, because that looks like what you need, only you have a Q(x) matrix for x'=(C)(x)+Q(x)

OpenStudy (anonymous):

I'm sorry but I have no idea to what you are referring to?

OpenStudy (anonymous):

My bad, what I meant was have you all learned to take that second order eq and turn it into a system?

OpenStudy (anonymous):

Where in that system, that second order becomes two first order equations and then you have them expressed in terms of matrix vectors.

OpenStudy (turingtest):

I just found my mistake in L's section, go check it out!

OpenStudy (anonymous):

ok will do! thanks so much!

OpenStudy (anonymous):

so solve the eq for y''\[y''=y+\sin(x)/7\] let \[ x_1=y; x_1'=y'=x_2; x_2'=y''\] so then y'' is \[x_2'=x_1+\sin(x)/7\] now you have two eqs \[x_1'=x_2\] \[x_2'=x_1+\sin(x)/7\]

OpenStudy (anonymous):

This becomes a matrix X'=k (x) + f(x) \[\left(\begin{matrix}x_1' \\ x_2'\end{matrix}\right)=\left[\begin{matrix}0 & 1 \\ 1 & 0\end{matrix}\right]\left(\begin{matrix}x_1 \\ x_2\end{matrix}\right)+\left(\begin{matrix}0 \\ \sin(x)/7\end{matrix}\right)\]

OpenStudy (turingtest):

@daomowon I already solved it without the matrix in the section L differential equations, just in case you're curious

OpenStudy (anonymous):

from there you treat the entire eq as if you would a homogeneous system of first order differential eqs. and solve for eigenvalues and vectors. Then take the Wronskian matrix of your eigenvectors and homogeneous solutions then solve the system with sin(x)/7 involved to get your actual equations by doing variation of parameters. and getting an alternate version of that first order system with respect to u'(x). It is way too hard to write the formula in with the subscripts, but it's reasonable when you look at it like that. Out of all the techniques I've learned, that is the only one I see possible. Good thing that you don't have a complete function matrix. Otherwise your last inegrals would be harder. Check your notes on variation of parameters dealing with systems of nonhomogeneous first order linear systems and that will get you through to a final solution.

OpenStudy (anonymous):

where can I see it?

OpenStudy (anonymous):

cool thanks